at a certain mountain location where the barometer reads 13.5 psia, the intake manifold gauge on a cross-country truck reads 20 inches of mercury vacuum. calculate the absolute pressure in the truck intake manifold, psia and in. of Hg.
1
Expert's answer
2020-03-23T12:23:30-0400

 The intake manifold gauge on a cross-country truck  reads 20 In Hg (inches Of Mercury) Vacuum (i.e. -20 In Hg Gauge).


"p=-20+13.5(2.03602)=7.5\text{ in. Hg}"

"p=-20(0.491154)+13.5=3.7\text{ psia}"